can you help me sirs...
2/5 of the students in a school were girls and the rest were boys.1/2 of the girls and 1/2 of the boys took part in the school sports.How many students were in the school if 570 students did not take part in the sports.

Answer by ptaylor(2198)

Let x=number of students in school
(2/5)x=number of girls
(1/2)(2/5)x==(1/5)x=number of girls who took part
Then (3/5)x=number of boys
And (1/2)(3/5)x=(3/10)x=number of boys who took part
So our equation to solve is:
(1/5)x+(3/10)x+570=x multiply each term by 10
2x+3x+5700=10x subtract 10x and also 5700 from each side
2x+3x-10x+5700-5700=10x-10x-5700 collect like terms
-5x=-5700 divide each side by -5
x=1140------------------------------------number of students in school
CK
(1/5)*1140+(3/10)*1140+570=1140
228+342+570=1140
1140-1140

Hope this helps---ptaylor